OpenStudy (anonymous):

Solve: y/-2 +5=13 A) -16 B) 16 C) -36 D) 36

OpenStudy (anonymous):

y/-2 + 5 = 13 y - 10 = -26 y = -16

OpenStudy (anonymous):

b

OpenStudy (anonymous):

\[\frac{ y }{ -2 }+5=13\]First multiply by -2 to get rid of the denominator on the bottom of y.\[-2(\frac{ y }{ -2 }+5=13)\]\[y-10=-26\]Then add 10 to both sides.\[y=-26+10\]\[y=-16\]

OpenStudy (anonymous):

^^ can u help me with a question